Charles recieved grades of 82, 93, 77, and 88 on his first four tests. What grade must he get on the fifth test to have an average of 89?
Answer by askmemath(368)

If 89 is the average then the total of 5 tests must be 89x5 = 445
Given the first four test scores 82, 93, 77, and 88 whose sum is =340
So number of marks required to get an average of 89 = 445-340 = 105
Another way of doing this problem is as follows
Since 89 is the average, add the difference between the average and the remaining numbers
So 82 - 89 = -7
93- 89 = +4
77-89 = -12
88-89 = -1
-7 + +4 + -12 + -1 = -16
Which means the final score is 16 more than 89
89+ 16 = 105